If two collinear vectors A and B are added, the resultant has a magnitude equal to 4.0. If B is subtracted from A, the resultant has a magnitude equal to 8.0. What is the magnitude of A?
1
Expert's answer
2020-12-24T11:26:52-0500

Vectors  A and B are collinear: "B=kA."

If two collinear vectors A and B are added


"|A+kA|=4"

If B is subtracted from A


"|A-kA|=8"

"\dfrac{|1-k|}{|1+k|}=\dfrac{8}{4}"

"|1-k|=2|1+k|, k\not=\pm1"

If "k>1"


"k-1=2+2k""k=-3, False"

If "-1< k<1"

"1-k=2+2k""k=-\frac{1}{3}, True"

If "k<-1"


"1-k=-2-2k""k=-3, True"

"|A||1+(-3)|=4"

"|A|=2"

The magnitude of A is 2.



"|A||1+(-\frac{1}{3})|=4"

"|A|=6"

The magnitude of A is 6.